OpenStudy (anonymous):

2cos^2 x + 3cosx + 1 = 0 Solve on the interval {0,2Pi} A. x= Pi/6 , x = 7pi/6 B. x= Pi, x = 2pi/3, x = 4pi/3 C. x = 2pi, x = pi/3 D. x = 2pi, x = pi/4, x= 5pi/4

OpenStudy (anonymous):

@countonme123

OpenStudy (anonymous):

you solve it like quadratic : let cosx = y then 2y^2 + 3y + 1 =0 2y^2 +2y + y + 1 =0 2y(y+1) + 1(y+1)=0 (2y+1)(y+1) =0 y = -1 , y = -1/2 so cos(x) = -1 , cos(x) = -1/2 can you answer now ?

OpenStudy (anonymous):

(another thing to remember is that those functions are periodic so for cos(x) = -1/2 there are some solutions for example)

OpenStudy (anonymous):

So it is c right?

OpenStudy (anonymous):

b.... nvm

OpenStudy (anonymous):

b